Woolworth wants to increase its total revenue from Zambia market. The company already identified two type of customers as high income and medium income and plans to offer a 10%


discount on the cloths it sells. The table shows the sells made from the two groups


High income customers


(sales per month in Kwacha)


Medium income customers


(sales per month in Kwacha)


Sales before 10% discount 1.55 million 1.50 million


Sales after 10% discount 1.65 million 1.70 million


(a) Calculate the price elasticities of demand for high income customers and medium income


customers.


(b) Do you support the discount applied for both type of customers? Explain based on your


results from (a)



1
Expert's answer
2022-02-23T12:22:26-0500

(a) Calculate the price elasticities of demand for high-income customers and medium-income customers.


Elasticity measures the responsiveness of one variable to changes in another variable.

The formula for price elasticity of demand is given by:


Price Elasticity of Demand ="\frac{percent\ change\ in\ quantity}{percent\ change\ in\ price}"


Percentage change in demand by high income is:

Percentage change in demand ="\frac{change\ in\ quantity\ demanded}{previous\ price}"

Percentage change in demand ="\frac{1.65 - 1.55}{1.55} = \frac{0.1}{1.55}=0.0645 = 6.45\%"


Price Elasticity of Demand of high income:

Price Elasticity of Demand ="\frac{percent\ change\ in\ quantity}{percent\ change\ in\ price}"


Price Elasticity of Demand ="\frac{6.45\%}{10\%}" = 0.65


Percentage change in demand by medium income is:

Percentage change in demand ="\frac{change\ in\ quantity\ demanded}{previous\ price}"

Percentage change in demand ="\frac{1.70 - 1.50}{1.50} = \frac{0.2}{1.50}=0.1333 = 13.33\%"


Price Elasticity of Demand of medium-income:

Price Elasticity of Demand ="\frac{percent\ change\ in\ quantity}{percent\ change\ in\ price}"


Price Elasticity of Demand ="\frac{13.33\%}{10\%}" = 1.33


(b) Do you support the discount applied for both types of customers? Explain based on your results from (a)

For high income:

Since the price elasticity of demand is below 1(0.65) in high-income clients, the demand is inelastic. The discount will not improve the company's profit but it will lead to a loss since it does not improve the sales overall. Therefore, I do not support the discount for this category.


For medium-income:

Because the price elasticity of demand is greater than 1 (1.33), the demand is elastic. If a discount is applied for this category, sales will increase and consequently profit will increase. Therefore, I support the application of a 10% discount for this category.